OpenStudy (anonymous):

the equatioj of the tangent to tje vurve y2/10 - x2/4 at the point 6,-10

OpenStudy (anonymous):

@dumbcow okey so i totualy forgot how to do these

OpenStudy (dumbcow):

ok equation for tangent line is: \[y = \frac{dy}{dx}(x - x_1) + y_1\] (x1,y1) is given point

OpenStudy (dumbcow):

does the equation = 1 by chance ? y^2/10 - x^2/4 = 1

OpenStudy (anonymous):

ohh sorry yea =1

OpenStudy (dumbcow):

to get dy/dx you need to take implicit derivative \[\rightarrow \frac{2y}{10} \frac{dy}{dx} - \frac{2x}{4} = 0\] solve for dy/dx \[\frac{dy}{dx} = \frac{(2x/4)}{(2y/10)} = \frac{x}{2}*\frac{5}{y} = \frac{5x}{2y}\]

OpenStudy (anonymous):

but doesnt make any difference right

OpenStudy (dumbcow):

final answer \[y = (\frac{5x_1}{2y_1})(x - x_1) + y_1\] plug in given point

OpenStudy (dumbcow):

what do you mean?

OpenStudy (anonymous):

hold on after getting dy/dx what did u do ?

OpenStudy (dumbcow):

substitute it into equation for tangent line